what is the height of a triangle with a base length of 4 and an area of 1620?
Found 2 solutions by jorel1380, nerdybill:
Answer by jorel1380(3719)

A=1/2bxh
1620=1/2x4xh
1620=2h
810=h..

Answer by nerdybill(7384)

what is the height of a triangle with a base length of 4 and an area of 1620?
.
since area of any triangle is:
(1/2)*base*height
.
(1/2)4h = 1620
4h = 3240
h = 810 (height)